Convert the following DMS values to Decimal Degrees or Decimal Degree values to DMS. Remember to...

  1. Convert the following DMS values to Decimal Degrees or Decimal Degree values to DMS. Remember to use all the appropriate symbols & directions (o,,, N, S, E, W)! -
    1. 50o22’ N = how many decimal degrees?
    2. 11o56’45” S = how many decimal degrees?
    3. 45o13’30”E = how many decimal degrees?
    4. 7o S = what reading in DMS?
    5. 14.3o W = what reading in DMS?
    6. 72.861o N = what reading in DMS?

  1. Measure the distance in kilometers between these two coordinates:

36o 55’ 20” N, 103o 55’ 33” W and 37o 4’ 36” N, 108o28.26” W
Answer #1

The symbols and conversions are elaborated below:

' = minute

^{\circ} = degree

" = second

DMS = Degrees Minutes Seconds

DD = Decimal Degrees

1 degree = 1 hour

1 minute = 60 seconds

1 hour = 60 minutes

DMS values can be converted to decimal degree values using the formula shown below:

min sec DD(Decimal Degrees) = ddegrees) + 60 + 3600 ...... (1)

Calculate the decimal degrees from the DMS values using equation (1):

(a)

min sec DD = d+ 60 + 3600 co

Therefore, the decimal degree value is 50,367N .

(b)

56 45 min sec DD= d+60 + 3600 * 60+ 3600 = -11.945s

Therefore, the decimal degree value is -11.9455 .

(c)

min sec 13 30 5+ 60 + 3600 = 45.2 +60 + 3600 = 45+ 2+

Therefore, the decimal degree value is 45.225E .

(d)

Calculate the DMS value from the given DD value (7 S) as shown below:

Degree = 7

Minutes = 0 X 60 = 0

Therefore, DMS reading is 7°00s .

(e)

Calculate the DMS value from the given DD value (14.3W) as shown below:

Degree = 14

Minutes = 0.3 x 60 = 18

Seconds = 0 X 60 = 0

Therefore, DMS reading is 14°180 W .

(f)

Calculate the DMS value from the given DD value (72.861N) as shown below:

Degree = 72

Minutes = 0.861 x 60 = 51.66

Seconds = 0.66 x 60 = 39,6

Therefore, DMS reading is 72°5140N .